OpenStudy (anonymous):

Which expression is the first step you take when finding the quotient of 5/6 (divide) 2/3?

OpenStudy (anonymous):

multiply by the reciprocal of \(\frac{2}{3}\) the reciprocal of \(\frac{2}{3}\) is \(\frac{3}{2}\)

OpenStudy (anonymous):

so first step is to write \[\frac{5}{6}\times \frac{3}{2}\]